Botox Facial Care had earnings after taxes of $282,000 in 20X1 with 200,000 shares of stock outstanding. The stock price was $81.80. In 20X2, eamings after taxes increased to S418,000 with the same 200,000 shares outstanding. The stock price was $94.00 a. Compute earnings per share and the P/E ratio for 20X1. (The P/E ratio equals the stock price divided by earnings per share.) (Do not round intermediate calculations. Round your final answers to 2 decimal places.) Earnings per share P/E ratio times b. Compute earnings per share and the P/E ratio for 20X2. (Do not round intermediate calculations. Round your final answers to 2 decimal places.) Earnings per share P/E ratio imes c. Why did the P/E ratio change? (Do not round intemediate calculations. Input your answers as percents rounded to 2 decimal places.) The stock price percent while EPS percent.
Answer #1

a) EPS and PE for 20X1

EPS = Net income available to equity shareholders / Shares Outstanding

= 282000/200000

= $1.41

P/E Ratio = Market Price per share / EPS

= 81.80/1.41

= 58.01 times

b) EPS and PE for 20X2

EPS = Net income available to equity shareholders / Shares Outstanding

= 418000/200000

= $2.09

P/E Ratio = Market Price per share / EPS

= 94/2.09

= 44.98 times

c) Why did P/E ration change

Change in stock price = (94-81.80)/81.80

= 12.2 /81.80

= 0.14914425427

= 14.91%

Change in EPS = (2.09-1.41)/1.41

= 0.68 /1.41

= 0.48226950354

= 48.23%

The stock price increases by 14.91% while EPS increases by 48.23%. The P/E ratio decreases as Change in EPS exceeds Change in stock price